OpenStudy (anonymous):

Write In 7=x in exponential form

OpenStudy (empty):

Any ideas?

OpenStudy (anonymous):

No sorry

OpenStudy (sleepyhead314):

Hello @wagnermar ! Sorry for the late reply :-/ \[\ln (x) = \log_{e}(x) = y \]\[when~~e^y = x\]so for example if I had ln 3 = x it would become e^x = 3 [just replace the 3 with your 7 to get what you want ;) ] I hope this helped! If you have any more questions, feel free to tag me @sleepyhead314 ^_^